# Engineering Spatial and Molecular Features from Cellular Niches to Inform Predictions of Inflammatory Bowel Disease

### v1

@@ 3.4.3 Feature importance analysis
Among the 44 input features, only the top 20 are shown in Figure 7. In the three-class case, niche–gene features dominated, led by niche_3_gene_MZT2A, followed by several Niche 4 genes (COL5A1, MZT2A, FZD1, CASP3). Other high-ranking features included niche_3_gene_HLA-DQB1, niche_1_gene_PIGR, and niche_4_gene_COL3A1, with few enrichment or composition variables (comp_niche_1, comp_niche_4) appearing. For the two-class task (HC vs. IBD), enrichment features dominated, especially enrichment_2-2 ($\mathrm{S}_{2,2}$), enrichment_3-1 ($\mathrm{S}_{3,1}$), and enrichment_2-3 ($\mathrm{S}_{2,3}$), while only a few niche–gene features ranked highly. Thus, separating UC from CD relied more on gene-level signals, whereas distinguishing HC from IBD depended on spatial interaction patterns.
@@ 4 Discussion
We sought to develop a MLP classifier capable of classifying IBD subtypes by engineering a multi-layered feature set from ST data. The model achieved $\sim 77\%$ accuracy in the more challenging three-class problem, with perfect classification of HC and lower accuracy between UC and CD. This is expected since distinguishing between UC and CD is challenging. A central finding of our work, which came out of the PI analysis, is the differential utility of our engineered features. PI analysis revealed that this task depended primarily on niche-gene features, highlighting subtype-specific molecular signals within these microenvironments. In comparison, the simpler two-class problem achieved $\sim 92\%$ accuracy, which was derived mainly from niche neighborhood enrichment scores, indicating that spatial cellular organization is a strong marker of general intestinal inflammation.
@@
The primary strength of our methodology lies in its novel, three-tiered feature engineering strategy, which captures niche composition, spatial enrichment, and localized gene expression to provide a holistic view of the tissue state. The use of NMF to define cellular niches provided an unbiased, data-driven approach to deconstructing complex spatial data into functional units. However, the small cohort of nine samples (three per condition) limits generalizability. The analysis was also constrained by the 980-gene panel; a whole-transcriptome approach could reveal more biomarkers. Future work must focus on validating these findings in larger, independent, and multi-center patient cohorts.
This framework has significant clinical and research implications. It offers a potential path toward a more objective, data-driven diagnostic tool to help resolve the ambiguity between UC and CD. The specific spatial patterns and molecular markers identified could form the basis of novel diagnostic assays. For research, our findings provide deeper insights into the distinct biological mechanisms driving these diseases. The causal graphs suggest testable hypotheses about how changes in niche composition and interactions drive pathology. Future studies could explore how these niche features evolve with therapy, potentially serving as predictive biomarkers for treatment outcomes, and should include functional investigations into the role of top-ranking genes in IBD pathogenesis.

### v2

@@ 3.4.3 Feature importance analysis
Among the 44 input features, only the top 20 are shown in Figure 8 (Appendix). In the three-class case, niche–gene features dominated, led by niche_3_gene_MZT2A, followed by several Niche 4 genes (COL5A1, MZT2A, FZD1, CASP3). Other high-ranking features included niche_3_gene_HLA-DQB1, niche_1_gene_PIGR, and niche_4_gene_COL3A1, with few enrichment or composition variables (comp_niche_1, comp_niche_4) appearing. For the two-class task (HC vs. IBD), enrichment features dominated, especially enrichment_2-2 ($\mathrm{S}_{2,2}$), enrichment_3-1 ($\mathrm{S}_{3,1}$), and enrichment_2-3 ($\mathrm{S}_{2,3}$), while only a few niche–gene features ranked highly. Thus, separating UC from CD relied more on gene-level signals, whereas distinguishing HC from IBD depended on spatial interaction patterns.
However, PI can be misleading under correlated or redundant features, especially for highly nonlinear learners (e.g., MLPs). In such cases, permuting a single correlated feature may over- or under-estimate its contribution due to shared information and interaction terms. We therefore interpret PI qualitatively and report it with its variability across permutations, noting this caveat.
@@ 4 Discussion
We sought to develop a MLP classifier capable of classifying IBD subtypes by engineering a multi-layered feature set from ST data. The model achieved $\sim 77\%$ accuracy in the more challenging three-class problem, with perfect classification of HC and lower accuracy between UC and CD. This is expected since distinguishing between UC and CD is challenging. A central finding of our work, revealed by PI analysis, is the differential utility of our engineered features. PI analysis revealed that this task depended primarily on niche-gene features, highlighting subtype-specific molecular signals within these microenvironments. In comparison, the simpler two-class problem achieved $\sim 92\%$ accuracy, which was derived mainly from niche neighborhood enrichment scores, indicating that spatial cellular organization is a strong marker of general intestinal inflammation.
@@
The primary strength of our methodology lies in its novel, three-tiered feature engineering strategy, which captures niche composition, spatial enrichment, and localized gene expression to provide a holistic view of the tissue state. The use of NMF to define cellular niches provided an unbiased, data-driven approach to deconstructing complex spatial data into functional units. However, the small cohort of nine samples (three per condition) limits generalizability. The analysis was also constrained by the 980-gene panel; a whole-transcriptome approach could reveal more biomarkers. In addition, we recognize the risk of overfitting from extensive hyperparameter tuning on small datasets [58]. To mitigate this, we employed subject-level group-stratified cross-validation, ensuring that FOVs from the same subject never appeared in both training and validation folds. We also used $L^{2}$ regularization and a compact feature space to constrain complexity. Nonetheless, given the limited number of subjects, the reported accuracies should be interpreted as cross-validated estimates within a proof-of-concept framework, pending external validation on larger, independent, and multi-center cohorts.
This framework has significant clinical and research implications. It offers a potential path toward a more objective, data-driven diagnostic tool to help resolve the ambiguity between UC and CD. The specific spatial patterns and molecular markers identified could form the basis of novel diagnostic assays. For research, our findings provide deeper insights into the distinct biological mechanisms driving these diseases. The causal graphs suggest testable hypotheses about how changes in niche composition and interactions drive pathology. However, these graphs can only model linear interactions. Future studies could explore how these niche features evolve with therapy, potentially serving as predictive biomarkers for treatment outcomes, and should include functional investigations into the role of top-ranking genes in IBD pathogenesis.
@@ 5 Conclusion
Appendix 0.A Appendix
Figure 7: Overview of the computational pipeline. Starting from NanoString CosMx spatial transcriptomics data (9 subjects, 171 FOVs, 459,095 cells), cell types were inferred using the cell2location algorithm, and NMF to identify four recurrent cellular niches. From these niches, 44 features were engineered describing composition, spatial organization, and gene expression. A MLP classifier trained using group-stratified cross-validation was then used to predict disease states (HC vs. UC vs. CD or HC vs. IBD). Model explainability analyses (causal discovery, statistical tests, and permutation importance) enabled biological interpretation of the compositional, spatial and molecular determinants of IBD.